# Scikit-fingerprints: easy and efficient computation of molecular fingerprints in Python

### v1

@@ 3.5 Complex pipelines for 3D fingerprints
For tasks requiring 3D information, i.e. fingerprints based on conformers, the whole processing pipeline becomes more complex. Conformers need to be generated and often post-processed with force field optimization, and resulting fingerprints may have missing values. Additionally, using more than one fingerprint is often beneficial, especially for virtual screening, as they take different geometry features into consideration. In Listing LABEL:code:3d_pipeline, we present an example how to create such pipeline for vectorizing molecules for screening, for GETAWAY [77] and WHIM [78] descriptors. This short example would require well over 100 lines of code in RDKit, even without parallelization.
@@ 3.6 Comparison with existing software
We compare our library to existing libraries for chemoinformatics, which also include molecular fingerprints computation. Differences are summarized in table 4.
We implement the largest number of fingerprints, including both all those available in other libraries, and new ones like MAP4 [79] or E3FP [80]. In terms of Python support, scikit-fingerprints is the first one to have a native Python package, with other libraries not supporting Python at all (CDK and OpenBabel), or just using an autogenerated wrapper (RDKit). We also fully support parallelism and even distributed computing, which is either nonexistent or very limited elsewhere. scikit-fingerprints is also the only library utilizing pre-commit hooks and dedicated security tools, and offering convenient, integrated datasets.
@@
Actively maintained? Yes No Yes Yes
@@
Official Python package No
No
Yes
Yes
@@
Easy commercial usage
@@
Table 4: Comparison of scikit-fingerprints with CDK and OpenBabel.
@@ 4 Impact
Lack of efficient parallelism is a major downside of existing solutions. Modern molecular databases can easily encompass millions of molecules, especially for virtual screening [11, 12]. Our solution, utilizing all available cores, results in significant speedups, enabling efficient processing of large datasets. This is also beneficial for hyperparameter tuning [27, 28], fingerprint concatenation [15], data fusion [29, 30], and other computationally complex tasks.
Simple class hierarchy and high code quality make our solution easily extensible. New fingerprints can easily be added, automatically benefiting from parallelization and scikit-learn compatibility. GitHub repository had 7 contributors to date, showing good reception by the community and easy learning curve. The first issue by an external researcher has been made in a week of making the library public, highlighting the need for modern software in this area.
The research shows that fingerprint-based molecular property prediction is still competitive compared to graph neural networks [13, 16, 14], justifying further research in this area. In particular, they should be applied as baselines for fair evaluation of the impact of novel approaches, which is particularly easy with our library. scikit-fingerprintshas already been applied to research in molecular chemistry. In [81], it was used to implement ECFP fingerprint as a baseline algorithm, ensuring fair comparison of various approaches on the MoleculeNet benchmark [56]. It is also being actively applied for predicting pesticide toxicity for honey bees, using recently proposed ApisTox dataset [82]. Additionally, numerous research projects and Master’s theses at Faculty of Computer Science at AGH University of Krakow are currently utilizing it.
Finally, scikit-fingerprintsis constantly evolving, with new fingerprints being added. We are also working on expanding the functionality, e.g. implementing data splitting functions based on fingerprints, or dataset loaders for popular benchmark datasets. Therefore, its impact in chemoinformatics will be even greater in the future.
@@ 5 Conclusions
We have developed scikit-fingerprints, an open-source Python library for computation of molecular fingerprints. It is simple to use, fully compatible with the scikit-learn API, and easily installable from PyPI. It is also the most feature-rich and highly efficient library available in the Python ecosystem, allowing parallel computation of over 30 different fingerprints. Multiple mechanisms have been implemented to ensure high code quality, maintainability, and security. It fills the gap for a single, definitive software in Python ecosystem for molecular fingerprints. It facilitates quicker, more efficient, and also more comprehensive experiments in fields of chemoinformatics, de novo drug design and computational molecular chemistry.
@@ Acknowledgements
Research was supported by the funds assigned by Polish Ministry of Science and Higher Education to AGH University of Krakow, and by the grant from “Excellence Initiative - Research University” (IDUB) for the AGH University of Krakow. We thank Michał Szafarczyk and Michał Stefanik for help with code implementation, and Wojciech Czech for help with manuscript review. We also thank Alexandra Elbakyan for her work and support for accessibility of science.

### v3

@@ 3.5 Complex pipelines for 3D fingerprints
For tasks requiring 3D information, i.e. fingerprints based on conformers, the whole processing pipeline becomes more complex. Conformers need to be generated and often post-processed with force field optimization, and resulting fingerprints may have missing values. Additionally, using more than one fingerprint is often beneficial, especially for virtual screening, as they take into account different geometry features. In Listing LABEL:code:3d_pipeline, we present an example of how to create such a pipeline to vectorize molecules for screening, calculating the GETAWAY [88] and WHIM [89] descriptors. This short example would require well over 100 lines of code in RDKit, even without parallelization.
@@ 3.6 Comparison with existing software
We compare scikit-fingerprints with existing libraries for chemoinformatics, which also support the computation of molecular fingerprints. Differences are summarized in Table 4.
In terms of Python support, we provide the first Python-native solution, with other libraries relying on various wrappers. It is also installable with pip from PyPI, and can be easily managed with modern dependency managers such as Poetry [90]. We implement the largest number of fingerprints, including both all those available in other libraries, and new ones like MAP4 [91] or E3FP [92]. Another advantage of scikit-fingerprints is the full support of parallelism and even distributed computing, which is nonexistent or very limited elsewhere. It is also the only library utilizing pre-commit hooks, dedicated security tools, and offering a fully scikit-learn compatible interface.
@@
pip-installable No Yes Yes Yes
Last PyPI update Never 2020 2024 2024
@@
scikit-learn compatible No No No Yes
@@
Easy proprietary usage
@@
Table 4: Comparison of scikit-fingerprints with other solutions.
@@ 4 Impact
The lack of efficient parallelism is a major downside of existing solutions. Modern molecular databases can easily encompass millions of molecules, especially for virtual screening [11, 12]. Our solution, utilizing all available cores, results in significant speedups, enabling efficient processing of large datasets. This is also beneficial for hyperparameter tuning [34, 35], fingerprint concatenation [15], data fusion [36, 37], and other computationally complex tasks.
Simple class hierarchy and high code quality make our solution easily extensible. New fingerprints can be easily added, automatically benefiting from parallelization and scikit-learn compatibility. GitHub repository had 7 contributors to date, showing a good reception by the community and an easy learning curve. The first issue by an external researcher has been made in a week of making the library public, highlighting the need for modern software in this area.
Research shows that fingerprint-based molecular property prediction remains competitive compared to graph neural networks [13, 16, 14], justifying further research in this area. In particular, they should be applied as baselines for a fair evaluation of the impact of novel approaches, which is particularly easy with our library. scikit-fingerprints has already been applied to molecular chemistry research. In [93], it was used to implement ECFP fingerprint as a baseline algorithm, ensuring fair comparison of various approaches on the MoleculeNet benchmark. It is also actively applied to predict the toxicity of pesticides for honey bees, using the recently proposed ApisTox dataset [94]. Furthermore, numerous research projects at the Faculty of Computer Science at AGH University of Krakow are currently utilizing it.
Finally, scikit-fingerprints is constantly evolving, with new fingerprints being added. We are also working on expanding the functionality, e.g. implementing data splitting functions based on fingerprints, or adding molecular filters like Lipinski’s Rule of 5 [95] for preprocessing. Therefore, its impact in chemoinformatics will be even greater in the future.
@@ 5 Conclusions
We have developed scikit-fingerprints, an open source Python library for computing molecular fingerprints. It is simple to use, fully compatible with the scikit-learn API, and easily installable from PyPI. It is also the most feature-rich and highly efficient library available in the open source Python ecosystem, allowing parallel computation of more than 30 different fingerprints. Multiple mechanisms have been implemented to ensure high code quality, maintainability, and security. It fills the gap for a single, definitive software in the Python ecosystem for molecular fingerprints. It facilitates quicker, more efficient, and more comprehensive experiments in the fields of chemoinformatics, drug design, and computational molecular chemistry.
@@ Acknowledgements
Research was supported by the funds assigned by Polish Ministry of Science and Higher Education to AGH University of Krakow, and by the grant from “Excellence Initiative - Research University” (IDUB) for the AGH University of Krakow. The authors thank Michał Szafarczyk and Michał Stefanik for their help with code implementation, Wojciech Czech for help with manuscript review, and Andrew Dalke for helpful discussions and comments. The authors also thank Aleksandra Elbakyan for her work and support for accessibility of science.